please solve the system
3x+4y=-2
x=-2y

Found 2 solutions by funmath, checkley71:
Answer by funmath(2873)

Solve the system
L1) 3x+4y=-2
L2) x=-2y
Substitute -2y in for x in L1) and solve for y:
3(-2y)+4y=-2
-6y+4y=-2
-2y=-2
-2y/-2=-2/-2
y=1
Substitute 1 in for y in L2):
x=-2(1)
x=-2
The solution is (x,y)=(-2,1)
Happy Calculating!!!

Answer by checkley71(8403)

X=-2Y
THEN
3(-2Y)+4Y=-2
-6Y+4Y=-2
-2Y=-2
Y=-2/-2
Y=1
X=-2Y
X=-2*1
X=-2
PROOF
3*-2+4*1=-2
-6+4=-2
-2=-2